OpenStudy (anonymous):

Flaws in a carpet tend to occur randomly and independently at a rate of one every 200 square feet. What is the probability that a carpet that is 8 feet by 10 feet contains no flaws?

OpenStudy (dumbcow):

area = 80 1/200 = 0.2/80 so mean is 0.2 flaws per 80 sq. ft \[P(k=0) = \frac{(0.2)^{0}e^{-.2}}{0!} = e^{-.2} = 0.8187\]

OpenStudy (anonymous):

the answer is 0.6703

OpenStudy (dumbcow):

haha oops...are we using poisson distribution here?

OpenStudy (anonymous):

yep

OpenStudy (dumbcow):

oh i had a stupid moment...can you spot my mistake? hint: 80 goes into 200 how many times

OpenStudy (anonymous):

2.5?

OpenStudy (dumbcow):

yeah...1/2.5 = 0.4 plug in 0.4 for mean instead, then it will work out :)

OpenStudy (anonymous):

Thank you.